# Geometrical Aspects Of Resources Distribution In Quantum Random Circuits

### v1

@@ B.1 Entanglement entropy
B.2 Multipartite systems with bipartite entanglement
@@
here, the $\rho_{i}=\mbox{Tr}_{i}|\psi\rangle\langle\psi|$ represent the reduced density matrix of the $i$-th qubit, obtained by tracing out the remaining $n_{q}-1$ qubits. The parameter $n_{q}$ denotes the total number of qubits. It follows that $0\leq Q\leq 1$. The distribution of $Q$ generated by a set of random unitary states, along with Haar measurements, is depicted in Fig. 13
Figure 13: Entanglement distribution for the states generated with our QRC function for eight qubits and different circuits depths. The image on the inset was taken from [7] and is included here for comparison.
B.3 Tangle
@@
B.4 Negativity
In tripartite systems, we can detect the presence of entanglement between subsystems by using the negativity $\mathcal{N}$, which is defined as follows [30]
Where $\rho^{tC}$ is the partial transpose of $\rho$ with respect to the subsystem C, and $\lambda_{i}\left(\rho^{tC}\right)$ are the eigenvalues of $\rho^{tC}$. The negativity can be equivalently interpreted as the sum of the absolute values of the negative eigenvalues of $\rho^{tC}$
B.5 Concurrence
B.6 Genuine measurement entanglement with concurrence

### v2

@@ A.1 Entanglement entropy
A.2 Negativity
It is possible to detect the presence of entanglement between subsystems using the negativity $\mathcal{N}$. In this case where we focus on tripartite systems, it is known as tripartite negativity, and is defined as follows [33]
where $\rho^{tC}$ is the partial transpose of $\rho$ with respect to the subsystem C, and $\lambda_{i}\left(\rho^{tC}\right)$ are the eigenvalues of $\rho^{tC}$. The negativity can be equivalently interpreted [34] as the sum of the absolute values of the negative eigenvalues of $\rho^{tC}$.
A.3 Concurrence
A.4 Multipartite systems with bipartite entanglement
@@
here, the $\rho_{i}=\mbox{Tr}_{i}|\psi\rangle\langle\psi|$ represent the reduced density matrix of the $i$-th qubit, obtained by tracing out the remaining $n_{q}-1$ qubits. The parameter $n_{q}$ denotes the total number of qubits. It follows that $0\leq Q\leq 1$. The distribution of $Q$ generated by a set of random unitary states, along with Haar measurements, is depicted in Fig. 18.
Figure 18: Entanglement distribution for the states generated with our QRC function for eight qubits and different circuit depths. The image on the inset was taken from [12] and is included here for comparison.
A.5 Tangle
@@
A.6 Genuine measurement entanglement with concurrence